OpenStudy (anonymous):

Write an equation in point slope form given: slope, m, = -1/5 and point (5, -2).

OpenStudy (anonymous):

\[-2-y _{1}=m(5-x _{1})\]

OpenStudy (anonymous):

A. y + 2 = -(1/5)(x - 5) B. y + 5 = -(1/5)(x - 2) C. y – 5 = (1/5)(x + 2) D. y – 5 = (1/5)(x - 2)

OpenStudy (texaschic101):

y - y1 = m(x - x1) slope(m) = -1/5 (5,-2) x1 = 5 and y1 = -2 now we sub y - 5 = -1/5(x - (-2) y - 5 = -1/5(x + 2) <=== point slope form

OpenStudy (texaschic101):

wait...I messed up,,,hold on

OpenStudy (texaschic101):

y - y1 = m(x - x1) y - (-2) = -1/5(x - 5) y + 2 = -1/5(x - 5)

OpenStudy (texaschic101):

sorry about that....I put the points in the wrong spot

OpenStudy (texaschic101):

answer is : y + 2 = -1/5(x - 5)

OpenStudy (texaschic101):

any questions ?